Im supposed to write each fraction as an equivalent fraction with the given denominator. Was doing good until they threw in letters. Can you help explain this process?

1/6 as ?/24x
Thanks

Answer by asha(30)

given 1/6 =?/24x
24x/6 =4x
therefore 1/6 =4x/24x
do not worry about the letters as you call them because they cancel when you divide the two terms.